OpenStudy (anonymous):

What is the range of the function f(x) = 3x − 12 for the domain {-2, 2}?

Nnesha (nnesha):

domain = x values range = y values so substitute x for given domain then solve for f(x) (which is same as y)

OpenStudy (anonymous):

@maddie82502

OpenStudy (anonymous):

Have you tried doing what Nnesha told?

OpenStudy (maddie82502):

@americansniper12 can u help

OpenStudy (anonymous):

The range is the y values. So you need to insert x=-2 and x=2 in your equation. Then you get the range.

OpenStudy (anonymous):

I ended up with (-6,-18)

OpenStudy (anonymous):

\[f(-2)=3*(-2)-12=?\] \[f(2)=3*2-12=?\]

OpenStudy (anonymous):

Yes, so that is your range, in the given interval :)